# Tiny models from tiny data: Textual and null-text inversion for few-shot distillation

### v1

@@ Appendix D Generated image examples
Figure 8: More examples of generated images, classes crate, guitar, roundworm, king crab, hourglass, dalmation, bowl and school bus. For each class, 5 random support examples are shown on top, followed by 10 randomly selected generated examples.
@@
Figure 8 shows more examples of images generated by TINT. The classes were selected manually, but the support examples and generated images were randomly selected without cherry picking. Note how for example the generated guitar images capture the setting of the support examples (focus on electric guitars / stage), rather than producing more generic guitar images. Also note how the generated king crab images cover both whole crabs and prepared as food, as well as covering both the red and brown color scales of the support examples. Sometimes food-like images are depicted in a brown hue, which is understandable based on the few support examples. The generator also sometimes extrapolates a bit too far in the food direction (lower-right image). The roundworm class is the one where the method struggles the most. Here, generated images are frequently reminiscent of atmospherical phenomena or pyrotechnics. This is understandable based on the two microscopy images present in the support set. Recall however that the negative effects of non-desired variation in the generated images are mitigated by the use of a teacher.

### v2

@@ Appendix C Generated image examples
Figure 9: More examples of generated images, classes crate, king crab, hourglass, dalmation, bowl, and school bus. For each class, 5 random support examples are shown on top, followed by 10 randomly selected generated examples.
@@
Figure 9 shows more examples of images generated by TINT. The classes were selected manually, but the support examples and generated images were randomly selected without cherry picking. Note how for example the generated guitar images capture the setting of the support examples (focus on electric guitars / stage), rather than producing more generic guitar images. Also note how the generated king crab images cover both whole crabs and prepared as food, as well as covering both the red and brown color scales of the support examples. Sometimes food-like images are depicted in a brown hue, which is understandable based on the few support examples. The generator also sometimes extrapolates a bit too far in the food direction (lower-right image).
Figure 10 shows a qualitative cross-domain example where the support examples are aerial images from the RSSCN dataset (Cheng et al., 2017). Although we expect that it is possible to improve the image quality by using a dedicated diffusion model, e.g. the model by Khanna et al. (2023), this example shows that the generated images stay reasonably faithful to the support examples even when a general diffusion model is used. As expected, the examples generated with low $\alpha$ values are rather similar to the support examples, while examples generated with higher $\alpha$ present more variation.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 10: Cross-domain example. Top row: support examples. Remaining rows: generated examples with $\alpha\in\{0.2,0.4,0.6,0.8\}$.
Appendix D Broader impact
This work has presented how to best train tiny classifiers to maximum accuracy using a minimum of application-specific data. We consider the work foundational in nature. As such, it is not tied to any direct application, but we acknowledge that it may potentially be used for applications with both positive and negative impact, ranging from medical applications to embedded military systems. Regarding the core image generation procedure, TINT, we acknowledge that generative models can be used to spread misinformation through deep fakes. In theory, it may be possible to use TINT to generate deep fakes by applying it to a set of images containing a certain individual, thereby generating similar-looking images of a similar-looking individual in new situations. However, this is a rather crude way of generating deep fakes compared to more direct image editing methods, and we do not expect our method to be adopted for this purpose.